# Analysis of reactor burnup simulation uncertainties for antineutrino spectrum prediction

### v2

@@ 3.5 Results
Table 4: Average Serpent C/E for chosen samples (SF95 2-5 and SF97 3-6) in the uniform temperature case (a) and with the model with temperature zones (b), compared with the literature results obtained with different simulation codes.
4 Fission fractions and $\bar{\nu}_{e}$ spectra results
Figure 8: Top: Fission fraction evolution in the SF95 rod predicted by simulation (a) with uniform temperature field and simulation (b) with non-uniform temperature field. Bottom: Relative fission fraction differences between simulation (b) and simulation (a) with respect to (b).
Figure 9: Top: Fission fraction evolution in the SF97 rod predicted by simulation (a) with uniform temperature field and simulation (b) with non-uniform temperature field. Bottom: Relative fission fraction differences between simulation (b) and simulation (a) with respect to (b).
4.1 Fission Fractions
In the top pads of Figures 8 and 9 we show the time evolution of the fission fractions in the Takahama-3 SF95 and SF97 rods, computed with Serpent simulations using uniform (a) or non-uniform (b) temperature field, as described in Sect. 3.4. With fresh fuel, the contribution to the fission events comes from ${}^{235}$ ($\sim$ 94%) and ${}^{238}$ ($\sim$ 6%). As the fuel depletion proceeds, two major effects can be appreciated: ${}^{235}$ is consumed and Pu nuclides are produced from ${}^{238}$ neutron captures. Therefore, with the increase in the burnup level, the relative importance to the total fission events of ${}^{235}$ tends to decrease, while the ${}^{239}$u and ${}^{241}$u contribution increases. The grey areas represent the cooling periods between cycles in which only radioactive decay occurs. In the bottom pads of Figures 8 and 9 we show the relative differences of fission fractions between simulation (a) and simulation (b). For both SF95 and SF97 rods, the relative differences have similar trends, with a weak dependence on burnup for ${}^{235}$, ${}^{238}$, ${}^{239}$u. For ${}^{241}$u we observe an unstable trend at the beginning of fuel depletion, which can be considered not relevant in our analysis since its fission fraction value is nearly zero at low burnup. It is worth noting here that on average, after the first operational cycle, the fission fractions predicted by simulation (b) differ from those predicted by simulation (a) by about:
  • (${}^{235}$) $+0.85\%$ in SF95 and $+0.23\%$ in SF97 rod;
  • (${}^{238}$) $-9.0\%$ in both SF95 and SF97 rods;
  • (${}^{239}$u) $+0.34\%$ in SF95 and $+0.92\%$ in SF97 rod;
  • (${}^{241}$u) $+1.1\%$ in SF95 and $+2.4\%$ in SF97 rod;
In light of these findings, we can state that modelling the temperature field in MC simulations has a non-negligible effect in the estimation of fission fractions, especially the ${}^{238}$ one. The temperature itself has a lower impact on the fissions from the ${}^{238}$. The relative reduction in the ${}^{238}$ fraction is addressed to the broadening of the fission cross sections resonances of the other nuclides (${}^{235}$, ${}^{239}$u, and ${}^{241}$u) in the hotter fuel volumes where the neutron flux is more intense. Thus, since the sum of fission fractions is normalized to 1, increasing the fission probability of the other three nuclides causes a decrease of the ${}^{238}$ fraction. Therefore, this analysis highlights that an overly approximate simulation model that does not take into account the temperature field can lead to fission fraction estimates affected by systematic errors on the scale of the differences listed above.
Finally, we also analyzed the fission fraction differences between the simulation (a), with respect to the one with a single burnup zone (simulation (c)). In this case the differences are $\lesssim 2\%$ for all nuclides.
4.2 Non oscillated $\bar{\nu}_{e}$ spectra results
Figure 10: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the fission fractions corresponding to three different burnup levels of SF97 rod. Bottom: Relative differences of spectra after 550 d and 1340 d with respect to the fresh fuel spectrum.
Figure 11: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the fission fractions at the end of the burnup cycle for the three simulations of the SF95 rod. Bottom: Relative differences of spectra from simulations (b) and (c) with respect to simulation (a).
We finally present the equilibrium $\bar{\nu}_{e}$ spectra obtained through the summation method by combining the cumulative fission yield (CFY) data published in the Live Chart of Nuclides by IAEA IAEA_livechart and the $\bar{\nu}_{e}$ spectra of 468 $\beta^{-}$ decays obtained through BetaShape by setting an energy binning of 1 keV. We use thermal CFY for ${}^{235}$, ${}^{239}$u and ${}^{241}$u, which are mostly fissioned by thermal neutrons, and fast CFY for ${}^{238}$ which can be fissioned with not negligible probability by fast neutrons only. Then, the total reactor $\bar{\nu}_{e}$ spectrum (Equation (2)) is evaluated with the fission fractions estimated with Serpent simulations. Since the reactor $\bar{\nu}_{e}$ will be detected through the IBD reaction, we multiply the $\bar{\nu}_{e}$ spectra by the IBD cross section STRUMIA200342 so as to obtain the shape of the detectable non oscillated spectrum. We want to highlight that this analysis is meant to compare the propagation of the reactor model differences on the reactor antineutrino spectrum. Other reactor antineutrino calculation methods as well as comparison with experimental data is out of the scope of the analysis and will be presented in future works.
In Fig. 10, we show the $\bar{\nu}_{e}$ spectra obtained by using the fission fractions estimated by Serpent for the Takahama-3 assembly at 3 different burnup levels for the SF97 rod: fresh fuel, intermediate burnup fuel (after 550 days), and end-of-life fuel (1340 days). As shown in the bottom pad, the $\bar{\nu}_{e}$ flux intensity decreases with time. This is because the average number of $\bar{\nu}_{e}$ emitted after ${}^{239}$u fission is lower with respect to that of ${}^{235}$. The relative difference calculated with respect to the fresh fuel $\bar{\nu}_{e}$ flux is of the order of 10% around the maximum of the spectrum at 4 MeV, and is up to $\sim 30\%$ in the high energy range. This result reflects the extreme case where the reactor starts from the fresh condition, without any refuelling strategy: in a more realistic scenario, this effects is averaged out considering fuel at different depletion stages. Nevertheless, in the framework of the JUNO experiment, it will be very important to correctly sum up the non-oscillated spectra taking into account the fission fraction evolution of each of the eight reactors during the data taking.
Finally, in Fig. 11 we show the impact of the systematic errors discussed in Sect. 4.1 on the non oscillated $\bar{\nu}_{e}$ spectrum. In particular, the spectra shown in the top pad were computed by using the fission fractions obtained from the simulations (a), (b) and (c) of the SF95 rod at the end of its second cycle (875 days). In this case, the differences are much less pronounced, remaining below 1% up to 7 MeV and in any case below 3% at higher energies. This can be addressed to the fact that the major differences are related to the ${}^{238}$ fraction, which has a minor impact on the overall system.
5 Conclusions
In this paper we presented a methodology for the prediction of the non-oscillated reactor $\bar{\nu}_{e}$ spectrum. This is intended to serve the scientific community in future reactor $\bar{\nu}_{e}$ experimental campaigns (e.g., JUNO and TAO) that will analyze the mass ordering problem with an unprecedented level of precision. We coupled the Monte Carlo code Serpent along with the $\bar{\nu}_{e}$ spectrum calculation software BetaShape. The Serpent simulations allow the calculation of fission fraction evolution over time, leveraging data encompassing reactor geometry, power history, and fresh fuel inventory. Subsequently, by integrating BetaShape alongside nuclear databases, we construct the antineutrino spectra emanating from the four primary fissile isotopes with the summation method – ${}^{235}$, ${}^{238}$, ${}^{239}$u, ${}^{241}$u – thereby enabling predictions of the total reactor antineutrino spectra across various burnup levels.
In this paper we focused our analysis on the fuel burnup, aiming at exploring the accuracy in the evaluation of fission fractions through MC simulations. For this purpose, we used the spent fuel concentrations of the Takahama-3 PWR as experimental benchmark for the Serpent simulations. We found that Serpent simulation run with either uniform temperatures in the fuel/moderator or imposing different temperature zones, are both in a reasonably good agreement with the experimental data. Importantly, our simulations highlight that the temperature field introduces minor discrepancies in the fuel depletion calculations, its effect but its influence is more pronounced when considering the fission fractions. Looking ahead, our future plans encompass harnessing this analytical tool to further investigate additional sources of systematic uncertainty (such as cross-section or fission yield data uncertainties) and their subsequent impact on the antineutrino spectrum. Future developments will be devoted to relax the equilibrium hypothesis on the fission products, highlighting the importance that this amount of nuclide have on the spectrum. Also, a more detailed analysis of the reactor configuration of the Taishan and Yangjiang plant will be addressed, along with the comparison with other available experimental spectra from near-detector experiments.

### v3

@@ Appendix A SF97 validation
Table 6: Average Serpent C/E for chosen samples (SF97 3-6) in the uniform temperature case and average boron (B), compared with the literature results obtained with different simulation codes.
Appendix B
Appendix B Other Fission fractions
Section 4.1 showed the comparison in fission fraction concerning the boron approach. This appendix aims at comparing the other important scenarios: one burnup zone at constant properties (A) and 30 burnup zones at variable temperature and constant boron (C).
Figure 14: Comparison between fission fraction in simulation A (blue line) and simulation B (green line) and their relative difference (red dashed line). The effect of spatial discretization for the burnup zones on the fission fraction calculation has a maximum in the first burnup cycle, and then tends to a uniform and negligible value in the second cycle.
The burnup discretization effect is showed in Figure 14: despite an initial discrepancy, around the burnup level where the gadolinium is depleted, the fission fraction difference tend to zero in the second cycle. From a simulation point of view, this helps in the simulation of large systems, where a fine discretization is not suitable.
Figure 15: Comparison between fission fraction in simulation B (green line) and simulation C (purple line) and their relative difference (red dashed line). The effect of spatial dependent temperature approach against the average treatment on the fission fraction calculation shows a negligible trend for all the isotopes, with a balance between 235U (-0.75 %) and 238U (+0.75 %).
The impact of the temperature field approach is shown in Figure 15: this comparison shows a slight dependence on burnup, decreasing the importance of 235U and increasing the 238U, laying always below the 1%. Plutonium fraction effects tend to zero in the second cycle, being unsensible to the temperature approach adopted. Also this effect, give useful information from a simulation point of view, allowing to treat the temperature as an average quantity, helping the description of large systems. Within the JUNO/TAO framework, such difference can be considered negligible.
Figure 16: Comparison between fission fraction in simulation B (green line) and simulation E (brown line) and their relative difference (red dashed line). The combined effect of the boron treatment and spatial temperature shows a similar behaviour with respect to the B-D comparison, the combined effect matches with a superimposition effect within the design choices.
Finally, it is worth mentioning that the combined effect follows a superimposition effect: Figure 16 compares the most simple case (B) with the most accurate one (E). Analyzing burnup data, the combined effect of temperature field and boron treatment is the sum of the single scenarios previously shown.